A vehicle has a wheel base of 3 m with a centre of gravity 0.8 m above ground level and 1.8 m in front of the rear axle. The coefficient of friction between the wheels and the road is 0.5. The vehicle is traveling horizontally at 90 km/h. Calculate the shortest distance at which the vehicle can be brought to a standstill by means of: 1.1. The brakes on the rear wheels only (14)

1.2. The brakes on both front and rear wheels. 


1
Expert's answer
2021-10-11T07:01:38-0400

L = 3 m ; x = 0.8 m ; h = 1.8 m ;

μ = 0.5

Let   a = Maximum deceleration of the car,

m = Mass of the car,

FA and FB = Braking forces at the front and rear wheels respectively, and

RA and RB = Normal reactions at the front and rear wheels respectively.


rakes on the rear wheels only:

We know that when the brakes are applied to all four wheels and the vehicle moves on a level road, then

"R _B=m⋅g( \frac{L−μ⋅h−x}{L}) = m⋅9.81( \frac{3−0.5⋅1.8−0.8}{3}) = 4.28m"


"F_{ A }+F_{ B }=m .a"


"2\u03bc.R _B=m\u22c5a"


"2\u00d70.5\u00d74.28m=m.a"


"a =4.28 \frac{m}{s^2}"